OpenStudy (anonymous):

7) If sin A+sin^2 A=1,find cos^2 + cos^4 A

myininaya (myininaya):

are you missing an A?

myininaya (myininaya):

zarkon!

OpenStudy (zarkon):

looks that way

OpenStudy (anonymous):

yea! missed one A..

OpenStudy (anonymous):

If sin A+sin^2 A=1,find cos^2 A + cos^4 A

myininaya (myininaya):

\[\sin^2(A)+\sin(A)-1=0\] solve \[x^2+x-1=0\] \[x=\frac{-1 \pm \sqrt{5}}{2}\] -------------- if \[\sin(A)=\frac{-1+\sqrt{5}}{2}\] Imagine we have a right triangle again sin means opp/hyp=(-1+sqrt{5}}/2 \[adj=\sqrt{2^2-(-1+\sqrt{5})^2}=\sqrt{4-(1-2\sqrt{5}+5)}=\sqrt{-2+2\sqrt{5}}\] so \[\cos(A)=\frac{-2+2\sqrt{5}}{2}=>\cos^2(A)=\frac{(-2+2\sqrt{5})^2}{2^2}\] and \[\cos^4(A)=\frac{(-2+2\sqrt{5})^2}{2^4}\]

myininaya (myininaya):

gosh i did something wrong

myininaya (myininaya):

sctrach everything i said after so cos(A)=

myininaya (myininaya):

\[\cos(A)=\frac{\sqrt{-2+2\sqrt{5}}}{2} \] there we go

myininaya (myininaya):

\[\cos^4(A)+\cos^2(A)=\frac{(\sqrt{-2+2\sqrt{5}})^4}{2^4}+\frac{(\sqrt{-2+2\sqrt{5}})^2}{2^2}\]

myininaya (myininaya):

\[=\frac{(-2+2\sqrt{5})^2+4(-2+2\sqrt{5})}{16}\]

myininaya (myininaya):

\[=\frac{4-8\sqrt{5}+20-8+8\sqrt{5}}{16}=\frac{-4}{16}=\frac{-1}{4}\]

myininaya (myininaya):

the answer turned out beautiful didnt expect that

myininaya (myininaya):

wait made a mistake

OpenStudy (zarkon):

you must have made a mistake it can't be negative

myininaya (myininaya):

\[=\frac{4+20-8}{16}=\frac{24-8}{16}=\frac{16}{16}=1\]

myininaya (myininaya):

i forgot the +20 there

myininaya (myininaya):

completely ignored him

OpenStudy (zarkon):

correct !

myininaya (myininaya):

but i probably did it the long way and usually when you choose long way you will make mistakes

OpenStudy (zarkon):

want to see another way ;)

myininaya (myininaya):

lol yes shorter?

myininaya (myininaya):

wait!

myininaya (myininaya):

let me try something else

OpenStudy (zarkon):

shorter...maybe not...though I don't need to solve for A

OpenStudy (zarkon):

or the sin(A)

myininaya (myininaya):

ok go ahead

OpenStudy (zarkon):

ok...I'm going to write in my latex editor then paste it here

myininaya (myininaya):

ok

OpenStudy (zarkon):

$$\cos^2(A)+\cos^4(A)$$ $$=\cos^2+(1-\sin^2(A))^2$$ $$=1-\sin^2(A)+1-2\sin^2(A)$$ $$=2-3\sin^2(A)+\sin^4(A)$$ $$=2-3\sin^2(A)+(\sin^2(A))^2$$ $$=2-3(1-\sin(A))+(1-\sin(A))^2$$ $$=2-3+3\sin(A)+1-2\sin(A)+\sin^2(A)$$ $$=\sin(A)+\sin^2(A)=1$$

OpenStudy (zarkon):

using the \[\sin^2(A)=1-\sin(A)\] which is given

OpenStudy (zarkon):

I think I might be missing a piece

myininaya (myininaya):

i see a mistake zarkon made a mistake :) well its only a little typo you are missing an A on 2nd line

OpenStudy (zarkon):

$$\cos^2(A)+\cos^4(A)$$ $$=\cos^2+(1-\sin^2(A))^2$$ $$=1-\sin^2(A)+1-2\sin^2(A)+\sin^4(A)$$ $$=2-3\sin^2(A)+\sin^4(A)$$ $$=2-3\sin^2(A)+(\sin^2(A))^2$$ $$=2-3(1-\sin(A))+(1-\sin(A))^2$$ $$=2-3+3\sin(A)+1-2\sin(A)+\sin^2(A)$$ $$=\sin(A)+\sin^2(A)=1$$

OpenStudy (zarkon):

lol that too...

OpenStudy (zarkon):

$$\cos^2(A)+\cos^4(A)$$ $$=\cos^2(A)+(1-\sin^2(A))^2$$ $$=1-\sin^2(A)+1-2\sin^2(A)+\sin^4(A)$$ $$=2-3\sin^2(A)+\sin^4(A)$$ $$=2-3\sin^2(A)+(\sin^2(A))^2$$ $$=2-3(1-\sin(A))+(1-\sin(A))^2$$ $$=2-3+3\sin(A)+1-2\sin(A)+\sin^2(A)$$ $$=\sin(A)+\sin^2(A)=1$$

OpenStudy (zarkon):

I believe I have it now

myininaya (myininaya):

yes looks good i see no mistakes

OpenStudy (zarkon):

good :)

myininaya (myininaya):

very nice zarkon

OpenStudy (zarkon):

you too

OpenStudy (anonymous):

wow! :D thanx! :D

myininaya (myininaya):

i keep on doing his the long way and you always find a shorter way for some reason i'm stuck on imagining a right triangle

OpenStudy (zarkon):

I first solved it on my calculator (doing it a similar way as you) ...so I knew the answer was 1 ;)

myininaya (myininaya):

i bet i will not change and i will keep doing it this way lol

myininaya (myininaya):

how do you do it on a cal?

myininaya (myininaya):

you said you have n-spire?

myininaya (myininaya):

or was that someone else?

OpenStudy (zarkon):

yes I had the calculator solve for A in \[sin(A)+\sin^2(A)=1\] for A (restricting A between 0 and 2pi) then evaluateed \[\cos^2(A)+\cos^4(A)\] with the A the calc found

myininaya (myininaya):

i probably can't do that with ti83

myininaya (myininaya):

well i guess i can graph the first one and approximate solution nvm i can do it

OpenStudy (zarkon):

you probably could...it just wouldn't be as nice...my nspire solved the first equation obtaining exact values...on the 83 you will get numerical approximations.

myininaya (myininaya):

i got 1.0000000071

myininaya (myininaya):

lol

myininaya (myininaya):

that was with approximation though

OpenStudy (zarkon):

I figured it would be a little off due to rounding

myininaya (myininaya):

so the answer is an approximation

OpenStudy (zarkon):

the nspire gave me the exact answer ... 1

myininaya (myininaya):

we can use newton's method to get us closer to this irrational number and then getting us closer to 1 lol

OpenStudy (anonymous):

what bout this sum guys?! If {x+cot40/tan50} - {1/2(cos35/sin35)}=4;find x

OpenStudy (zarkon):

just isolate x

myininaya (myininaya):

\[x=4-\frac{\cot(40)}{\tan(50)}+\frac{1}{2}*\frac{\cos(35)}{\sin(35)}\]

myininaya (myininaya):

just like solving x+3-5=4 for x x=4-3+5

myininaya (myininaya):

trig(some number) is just a regular number like any other number

OpenStudy (anonymous):

what bout this? P.T {sin theta*cos(90-theta)*cos theta/sin(90-theta)} + {cos theta*sin(90-theta)*sin theta/cos(90-theta)}